# Multimodal Adversarial Defense for Vision-Language Models by Leveraging One-To-Many Relationships

### v2

@@ Abstract
Pre-trained vision-language (VL) models are highly vulnerable to adversarial attacks. However, existing defense methods primarily focus on image classification, overlooking two key aspects of VL tasks: multimodal attacks, where both image and text can be perturbed, and the one-to-many relationship of images and texts, where a single image can correspond to multiple textual descriptions and vice versa (1:N and N:1). This work is the first to explore defense strategies against multimodal attacks in VL tasks, whereas prior VL defense methods focus on vision robustness. We propose multimodal adversarial training (MAT), which incorporates adversarial perturbations in both image and text modalities during training, significantly outperforming existing unimodal defenses. Furthermore, we discover that MAT is limited by deterministic one-to-one (1:1) image-text pairs in VL training data. To address this, we conduct a comprehensive study on leveraging one-to-many relationships to enhance robustness, investigating diverse augmentation techniques. Our analysis shows that, for a more effective defense, augmented image-text pairs should be well-aligned, diverse, yet avoid distribution shift—conditions overlooked by prior research. Our experiments show that MAT can effectively be applied to different VL models and tasks to improve adversarial robustness, outperforming previous efforts. Our code will be made public upon acceptance.
Vision-Language Model, Adversarial Robustness, Adversarial Defense
@@ 1 Introduction
Figure 1: Comparison of adversarial training (AT) methods for robust VL models. (a) Unimodal AT, such as TeCoA (Mao et al., 2022) and FARE (Schlarmann et al., 2024), robustifies a single modality via unimodal adversarial examples (AEs). However, it overlooks two key aspects of VL tasks: multimodal attacks, where attackers perturb both modalities, and one-to-many cross-modal alignment, where an image has multiple valid descriptions, and vice versa. (b) MAT addresses multimodal attacks by generating multimodal AEs during AT. (c) MAT+ further captures the inherent ambiguity in image-text relationships via one-to-many augmentations.
@@
Vision-language (VL) tasks require modeling the relationships between image and text data. For example, image-text retrieval (ITR) aims to retrieve the most relevant text given an image query, and vice versa. Recent approaches leverage VL models pre-trained on large-scale image-text pairs, such as CLIP (Radford et al., 2021), which learns joint representations for images and texts via contrastive learning. However, recent studies revealed that these models are vulnerable to adversarial attacks (Zhang et al., 2022; Lu et al., 2023), which exploit nearly imperceptible perturbations in the input data. This poses significant risks of causing unintended consequences in real-world applications. For instance, in recommendation systems, attackers may manipulate images or their descriptions to alter the retrieval rankings, unfairly promoting or demoting specific entities. As VL models are deployed into society, understanding and mitigating their adversarial vulnerabilities has become increasingly urgent.
@@
However, existing defense strategies for VL models (Mao et al., 2022; Wang et al., 2024b; Schlarmann et al., 2024) primarily focus on image attacks, e.g., for robust zero-shot image classification, leaving other VL tasks unexplored. This is a considerable oversight due to two critical aspects: (1) Multimodal manipulation: Attackers can manipulate both image and text modalities in VL tasks, while existing defense methods only consider image perturbations. Such an increased attack capability in VL tasks requires more complex defense strategies. (2) One-to-many (1:N) cross-modal alignment: Unlike classification, where text labels are simple and unambiguous (e.g., “a photo of a dog”), VL tasks involve sentence-level texts with a larger variation and ambiguity (e.g., the same image can be described as “a man with glasses is wearing a beer can crocheted hat” or “a man wears an orange hat and glasses”). This property makes achieving robust image-text alignment more complicated. Existing defense methods overlook these two key aspects, casting doubt on their effectiveness for VL tasks beyond zero-shot image classification.
@@
To address this gap, we pioneer a study on defense strategies for VL models against multimodal attacks. Specifically, we study how to robustly fine-tune VL models for downstream VL tasks. First, we propose multimodal adversarial training (MAT), which incorporates both image-text perturbations during training. MAT significantly enhances robustness against multimodal attacks, demonstrating the necessity of defense methods tailored for multimodal attacks, an aspect overlooked by vision-only unimodal strategies (Mao et al., 2022; Schlarmann et al., 2024).
@@
Furthermore, we argue that MAT’s performance is limited by the inadequate approximation of the real data distribution when using the deterministic (1:1) image-text pairs contained in the training data. To mitigate this, we investigate how to leverage the one-to-many (1:N) relationships inherent in VL data to enhance robustness. Inspired by works in cross-modal ambiguity modeling in ITR (Kim et al., 2023; Song & Soleymani, 2019), we explore various augmentation techniques to create diverse one-to-many (1:N) and many-to-one (N:1) image-text pairs. Our in-depth analysis reveals that augmentations are effective when image-text pairs are well-aligned and diverse, without causing distribution shift. For instance, text augmentations outperform image augmentations, whose higher dimensionality makes it difficult to avoid distribution shift. Additionally, cross-modal augmentations (e.g., $image\rightarrow text$) outperform intra-modal augmentations (e.g., $text\rightarrow text$) due to generating well-aligned image-text pairs. These are novel insights in the field of multimodal robustness, and complement the literature on unimodal adversarial training.
@@
  • First defense strategy against multimodal attacks in VL models: We demonstrate that existing defense methods focused on vision are suboptimal for robust VL tasks and pioneer research in this new direction.
  • Proposed multimodal adversarial training (MAT): We largely improve robustness against multimodal attacks, highlighting the importance of considering multimodal perturbations in VL data.
@@ 2 Related work
Adversarial attacks on vision-language models. Adversarial attacks on VL models are categorized into unimodal and multimodal. Unimodal attacks, such as gradient-based image attacks (Madry et al., 2017) and BERT-Attack for text (Li et al., 2020), perturb a single modality to mislead the models. In contrast, multimodal attacks, which perturb both image and text modalities, are significantly more effective (Zhang et al., 2022; Lu et al., 2023; Han et al., 2023; Wang et al., 2024a). However, developing defense strategies against multimodal attacks for VL tasks remains largely unexplored.
Adversarial defense for vision-language models. Existing defense strategies for VL models mainly focus on vision robustness, in which adversarial attacks perturb only the image. For example, Mao et al. (2022) and Wang et al. (2024b) approached zero-shot image classification on CLIP by proposing robust fine-tuning methods, which leverage unimodal adversarial training schemes to improve robustness. Schlarmann et al. (2024) also focused on image attacks only, and proposed a method for fine-tuning CLIP’s vision encoder to improve robustness in several VL tasks (i.e., image classification, image-text retrieval). Unlike the previous work, ours is the first to investigate adversarial defense strategies against multimodal attacks, where both image and text modalities are perturbed. Specifically, we propose a multimodal adversarial training strategy to enhance robustness against such attacks in VL models and tasks. Furthermore, we leverage the one-to-many (1:N) image-text relationships to further improve adversarial robustness.
Leveraging the one-to-many (1:N) nature of image-text. To tackle robustness in VL models, we take inspiration from current works that aim to model the ambiguity between image and text pairs; although a sentence can have multiple visual interpretations, normally only one is paired as the ground truth. Similarly, an image can be described using multiple different captions, but only one is considered as its pair. Since such 1:1 deterministic relationship is inconsistent with the 1:N nature of the data, existing works propose representing image-text samples as probabilistic embeddings (Chun et al., 2021; Chun, 2024), considering neighboring samples in the triplet loss (Thomas & Kovashka, 2020), and generating multiple and diverse representations for each image-text sample (Song & Soleymani, 2019; Kim et al., 2023). Among these solutions, the latter naturally fits the data augmentation strategy of adversarial training.
@@ 3 Preliminaries
Vision-language models. Many recent VL models (e.g., ALBEF (Li et al., 2021) and BLIP (Li et al., 2023)) are fundamentally built on CLIP, which learns joint image-text representations via large-scale image-text contrastive learning. CLIP consists of an image encoder $f_{\theta_{\text{I}}}:\mathbb{R}^{d_{\text{I}}}\to\mathbb{R}^{d_{\text{E}}}$ and a text encoder $f_{\theta_{\text{T}}}:\mathbb{R}^{d_{\text{T}}}\to\mathbb{R}^{d_{E}}$, where $\theta_{\text{I}}$ and $\theta_{\text{T}}$ are their parameters, $d_{\text{I}}$ and $d_{\text{T}}$ are input dimensions, and $d_{\text{E}}$ is the joint embedding dimension. Given an image $I\in\mathbb{R}^{d_{\text{I}}}$ and a text $T\in\mathbb{R}^{d_{\text{T}}}$, CLIP is trained to map them into a shared embedding space, maximizing the cosine similarity of image-text embeddings $S_{\theta_{\text{I},\text{T}}}(I,T)=\cos(f_{\theta_{\text{I}}}(I),f_{\theta_{\text{T}}}(T))$ for correct image-text pairs while minimizing it for incorrect pairs. CLIP optimizes the InfoNCE loss for a batch of $N$ image-text pairs $\{(I_{i},T_{i})\}_{i=1}^{N}$ as:
@@
$(I^{\prime},T^{\prime})$ $=(I+\delta_{\text{I}},T+\delta_{\text{T}})$ (2)
$=\operatorname*{arg\,min}_{\delta_{\text{I}}\in\Delta_{\text{I}},\delta_{\text{T}}\in\Delta_{\text{T}}}S_{\theta_{\text{I},\text{T}}}(I+\delta_{\text{I}},T+\delta_{\text{T}}),$
where $\delta_{\text{I}}$ and $\delta_{\text{T}}$ are image and text perturbations, and $\Delta_{\text{I}}$ and $\Delta_{\text{T}}$ define the set of allowed image and text perturbations. Image attacks maintain perceptual similarity between $I$ and $I^{\prime}$ typically using the $L_{p}$-norm. A common image attack strategy is projected gradient descent (PGD) (Madry et al., 2017), which iteratively updates $I^{\prime}$ by taking a small step in the direction of the gradient. Text attacks, such as BERT-Attack (Li et al., 2020), modify $N$ tokens in the text $T$ to maximize the divergence between $f_{\theta_{\text{T}}}(T)$ and $f_{\theta_{\text{T}}}(T^{\prime})$. Multimodal attacks perturb both image-text modalities to create $(I^{\prime},T^{\prime})$; Co-Attack (Zhang et al., 2022) perturbs each modality sequentially, first perturbing the image and then the text; SGA (Lu et al., 2023) enhances Co-Attack by considering the set-level interaction between multiple images and texts.
Adversarial training for image classification. Adversarial training (AT) (Madry et al., 2017) is a widely used defense against adversarial attacks, where a model is trained on adversarial examples. It solves a min-max optimization problem:
@@
Based on this, to improve CLIP’s adversarial robustness in zero-shot image classification, TeCoA (Mao et al., 2022) adversarially fine-tunes CLIP’s image encoder by minimizing the contrastive loss between adversarial images and the text expression of the corresponding class (“a photo of {*}”):
@@
However, TeCoA only defends against image adversarial attacks and does not consider the one-to-many (1:N) relationship in VL tasks.
@@ 4.1 Problem setup
To defend against multimodal attacks, we introduce MAT, which perturbs both image and text modalities during AT. Following the theory of unimodal AT (Eq. 3), we formulate MAT objective as:
@@ 4.2.1 Multimodal inner maximization
Directly optimizing the inner-maximization problem in Eq. 5 is highly complex, as it requires perturbing both modalities while accounting for their interactions. Nevertheless, in adversarial training, precise inner maximization is not always necessary; in fact, the basic PGD-AT defense method (Madry et al., 2017) uses approximation by training against first-order adversaries. Ensuring efficient optimization is equally important for feasible training time and scalability.
To this end, we propose two practical strategies for approximating the inner maximization: (1) Step-by-step perturbation: we generate adversarial examples $(I^{\prime},T^{\prime})$ by sequentially perturbing the text and image modalities, and (2) Loss function simplification: we replace the original loss function $\mathcal{L}_{\text{VL}}$ with a more efficient alternative when needed.
Adversarial text generation. First, we generate adversarial texts using a representative text attack, BERT-attack (Li et al., 2020). BERT-attack identifies critical words contributing to the loss and replaces them with semantically similar, grammatically correct alternatives to maximize the loss. However, this discrete optimization process is computationally expensive, and directly maximizing $\mathcal{L}_{\text{VL}}$ over multiple image-text pairs in a batch by repeatedly replacing words would be prohibitively time-consuming. To address this, we exploit the shared image-text alignment objective in VL models (e.g., CLIP, ALBEF, BLIP), and instead of direct loss maximization, we maximize the divergence between individual image-text embeddings as:
$T^{\prime}=\operatorname*{arg\,max}_{\delta_{\text{T}}\in\Delta_{\text{T}}}\|f_{\theta_{\text{T}}}(T+\delta_{\text{T}})-f_{\theta_{\text{I}}}(I)\|.$ (6)
Adversarial image generation. Next, given adversarial texts $T^{\prime}$, we generate adversarial images $I^{\prime}$ using PGD attack (Madry et al., 2017). Unlike the discrete optimization in text attacks, image attacks involve continuous optimization, where each update step requires a single back-propagation to maximize a loss function. Therefore, directly maximizing the loss $\mathcal{L}_{\text{VL}}$ is practically feasible. Nevertheless, for CLIP, instead of maximizing $\mathcal{L}_{\text{CLIP}}$, we adopt a more efficient approach by minimizing the cosine similarity between the image-text embeddings of $I^{\prime}$ and $T^{\prime}$ as:
$I^{\prime}=\operatorname*{arg\,max}_{\delta_{\text{I}}\in\Delta_{\text{I}}}-\frac{f_{\theta_{\text{I}}}(I+\delta_{\text{I}})\cdot f_{\theta_{\text{T}}}(T^{\prime})}{\|f_{\theta_{\text{I}}}(I+\delta_{\text{I}})\|\|f_{\theta_{\text{T}}}(T^{\prime})\|}.$ (7)

### v3

@@ Abstract
Pre-trained vision-language (VL) models are highly vulnerable to adversarial attacks. However, existing defense methods primarily focus on image classification, overlooking two key aspects of VL tasks: multimodal attacks, where both image and text can be perturbed, and the one-to-many relationship of images and texts, where a single image can correspond to multiple textual descriptions and vice versa (1:N and N:1). This work is the first to explore defense strategies against multimodal attacks in VL tasks, whereas prior VL defense methods focus on vision robustness. We propose multimodal adversarial training (MAT), which incorporates adversarial perturbations in both image and text modalities during training, significantly outperforming existing unimodal defenses. Furthermore, we discover that MAT is limited by deterministic one-to-one (1:1) image-text pairs in VL training data. To address this, we conduct a comprehensive study on leveraging one-to-many relationships to enhance robustness, investigating diverse augmentation techniques. Our analysis shows that, for a more effective defense, augmented image-text pairs should be well-aligned, diverse, yet avoid distribution shift—conditions overlooked by prior research. This work pioneers defense strategies against multimodal attacks, providing insights for building robust VLMs from both optimization and data perspectives.
@@ 1 Introduction
Figure 1: Comparison of adversarial training (AT) methods for robust VL models. (a) Unimodal AT, such as TeCoA [mao2022understanding] and FARE [schlarmann2024robust], robustifies a single modality via unimodal adversarial examples (AEs). However, it overlooks two key aspects of VL tasks: multimodal attacks, where attackers perturb both modalities, and one-to-many cross-modal alignment, where an image has multiple valid descriptions, and vice versa. (b) MAT addresses multimodal attacks by generating multimodal AEs during AT. (c) MAT+ further captures the inherent ambiguity in image-text relationships via one-to-many augmentations.
@@
Vision-language (VL) tasks require modeling the relationships between images and texts; for example, image-text retrieval (ITR) retrieves the most relevant text given an image query, and vice versa. Recent VL models such as CLIP [radford2021learning] achieve strong performance in these tasks; however, recent studies revealed that they are vulnerable to adversarial attacks [zhang2022towards, lu2023set], which exploit nearly imperceptible input perturbations. Such vulnerabilities pose serious practical risks; for instance, attackers may manipulate images or descriptions to alter retrieval rankings, unfairly promoting or demoting entities in recommendation systems. As VL models are increasingly deployed, understanding and mitigating their adversarial vulnerabilities is urgent.
@@
However, existing defense strategies for VL models [mao2022understanding, wang2024pre, schlarmann2024robust] primarily focus on image attacks (e.g., robust zero-shot image classification), leaving other VL tasks unexplored. This is a considerable oversight for two reasons: (1) Multimodal manipulation: Attackers can perturb both images and texts, requiring more complex defense strategies than image-only methods. (2) One-to-many (1:N) cross-modal alignment: Unlike classification with simple and deterministic labels (e.g., “a photo of a {class}”), VL tasks involve diverse, ambiguous sentences (e.g., “a man with glasses” vs. “a man wears an orange hat and glasses”), making robust image-text alignment more complicated. By overlooking these aspects, existing defenses are limited in their effectiveness beyond zero-shot image classification.
To address this gap, we pioneer defense strategies for VL models against multimodal attacks. Specifically, we study how to robustly fine-tune VL models for downstream VL tasks from both optimization and data perspectives, through extensive analysis.
@@
First, we propose Multimodal Adversarial Training (MAT), which incorporates both image-text perturbations during training. Perturbing both modalities is non-trivial due to (1) the difficulty of simultaneous updates, (2) multiple objective choices for image-text attacks, and (3) increased computational complexity. Through extensive analysis, we designed MAT to be both effective and reasonably efficient, while also offering valuable insights for future work. MAT significantly enhances multimodal robustness, demonstrating the necessity of defense methods tailored for multimodal attacks, an aspect overlooked by vision-only unimodal strategies [mao2022understanding, schlarmann2024robust].
@@
Furthermore, we argue that MAT’s performance is limited by the inadequate approximation of the real data distribution when using the deterministic (1:1) image-text pairs contained in the training data. To address this, we leverage the inherent one-to-many (1:N) relationships in VL data to enhance robustness. Inspired by works in cross-modal ambiguity modeling in ITR [kim2023improving, song2019polysemous], we explore augmentation techniques to create diverse one-to-many (1:N) and many-to-one (N:1) image-text pairs. Our in-depth analysis reveals that augmentations are effective when pairs are well-aligned and diverse, without inducing distribution shift. Specifically, text augmentations outperform image augmentations, since the higher dimensionality of images makes distribution shift harder to avoid. Moreover, cross-modal augmentations (e.g., $image\rightarrow text$) outperform intra-modal ones (e.g., $text\rightarrow text$) by generating better-aligned pairs. These findings provide novel insights into multimodal robustness and complement the literature on unimodal adversarial training.
@@
  • First defense strategy against multimodal attacks in VL models: We show that existing image-only defense methods are suboptimal for robust VL tasks and pioneer research in this new direction. Specifically, we investigate strategies from both optimization and data perspective.
  • Proposed Multimodal Adversarial Training (MAT): We designed MAT to be both effective and efficient, through extensive analysis. MAT largely improves multimodal robustness, highlighting the importance of considering multimodal perturbations in VL data.
@@ 2 Related work
Adversarial attacks on vision-language models. Adversarial attacks on VL models are categorized into unimodal and multimodal. Unimodal attacks, such as gradient-based image attacks [madry2017towards] and BERT-Attack for text [li2020bert], perturb a single modality to mislead the models. In contrast, multimodal attacks, which perturb both image and text modalities, are significantly more effective [zhang2022towards, lu2023set, han2023ot, wang2024transferable]. However, developing defense strategies against multimodal attacks for VL tasks remains largely unexplored.
Adversarial defense for vision-language models. Existing defense strategies for VL models mainly focus on vision robustness, in which adversarial attacks perturb only the image. For example, mao2022understanding and wang2024pre approached zero-shot image classification on CLIP by proposing robust fine-tuning methods, which leverage unimodal adversarial training schemes to improve robustness. schlarmann2024robust also focused on image attacks only, and proposed a method for fine-tuning CLIP’s vision encoder to improve robustness in several VL tasks (i.e., image classification, image-text retrieval). Unlike the previous work, ours is the first to investigate adversarial defense strategies against multimodal attacks. Specifically, we propose a multimodal adversarial training strategy to enhance robustness against such attacks in VL models and tasks. Furthermore, we leverage the one-to-many (1:N) image-text relationships to further improve adversarial robustness.
Leveraging the one-to-many (1:N) nature of image-text. Recent works aimed at modeling the ambiguity between image and text pairs; a sentence may have multiple visual interpretations and an image may be described in various ways, however, typically only one pair is used as ground truth. Such deterministic 1:1 pairing is inconsistent with the natural 1:N relationships in the data. To address this, prior studies propose representing image-text samples as probabilistic embeddings [chun2021probabilistic, chun2024improved], incorporating neighboring samples in the triplet loss [thomas2020preserving], and generating multiple diverse representations for each pair [song2019polysemous, kim2023improving]. Inspired by these works, we leverage 1:N relationships to augment data in adversarial training and enhance VL robustness.
@@ 3 Preliminaries
Vision-language models. Many recent VL models (e.g., ALBEF [li2021align] and BLIP [li2023blip]) are fundamentally built on CLIP, which learns joint image-text representations via large-scale image-text contrastive learning. CLIP consists of an image encoder $f_{\theta_{\text{I}}}:\mathbb{R}^{d_{\text{I}}}\to\mathbb{R}^{d_{\text{E}}}$ and a text encoder $f_{\theta_{\text{T}}}:\mathbb{R}^{d_{\text{T}}}\to\mathbb{R}^{d_{E}}$, where $\theta_{\text{I}}$ and $\theta_{\text{T}}$ are their parameters, $d_{\text{I}}$ and $d_{\text{T}}$ are input dimensions, and $d_{\text{E}}$ is the joint embedding dimension. Given an image $I\in\mathbb{R}^{d_{\text{I}}}$ and a text $T\in\mathbb{R}^{d_{\text{T}}}$, CLIP is trained to map them into a shared embedding space, maximizing the cosine similarity of image-text embeddings $S_{\theta_{\text{I},\text{T}}}(I,T)=\cos(f_{\theta_{\text{I}}}(I),f_{\theta_{\text{T}}}(T))$ for correct image-text pairs while minimizing it for incorrect pairs. CLIP optimizes the InfoNCE loss for a batch of $N$ image-text pairs $\{(I_{i},T_{i})\}_{i=1}^{N}$ as:
@@
$(\delta_{I}^{\star},\delta_{T}^{\star})$ $=\operatorname*{arg\,min}_{\delta_{I}\in\Delta_{I},\,\delta_{T}\in\Delta_{T}}S_{\theta_{I,T}}\big(I+\delta_{I},\,T+\delta_{T}\big),$ (2)
$(I^{\prime},T^{\prime})$ $=\big(I+\delta_{I}^{\star},\,T+\delta_{T}^{\star}\big).$ (3)
where $\delta_{\text{I}}$ and $\delta_{\text{T}}$ are image and text perturbations, and $\Delta_{\text{I}}$ and $\Delta_{\text{T}}$ define the set of allowed image and text perturbations. Image attacks maintain perceptual similarity between $I$ and $I^{\prime}$ typically using the $L_{p}$-norm. A common image attack strategy is projected gradient descent (PGD) [madry2017towards], which iteratively updates $I^{\prime}$ by taking a small step in the direction of the gradient. Text attacks, such as BERT-Attack [li2020bert], modify $N$ tokens in the text $T$ to maximize the divergence between $f_{\theta_{\text{T}}}(T)$ and $f_{\theta_{\text{T}}}(T^{\prime})$. Multimodal attacks perturb both image-text modalities to create $(I^{\prime},T^{\prime})$; Co-Attack [zhang2022towards] perturbs each modality sequentially, first perturbing the image and then the text; SGA [lu2023set] enhances Co-Attack by considering the set-level interaction between multiple images and texts.
Adversarial training for image classification. Adversarial training (AT) [madry2017towards] is a widely used defense against adversarial attacks, where a model is trained on adversarial examples. It solves a min-max optimization problem:
@@
Based on this, to improve CLIP’s adversarial robustness in zero-shot image classification, TeCoA [mao2022understanding] adversarially fine-tunes CLIP’s image encoder by minimizing the contrastive loss between adversarial images and the text expression of the corresponding class (“a photo of {*}”):
@@
However, TeCoA only defends against image attacks.
@@ 4.1 Problem setup
To defend against multimodal attacks, we introduce MAT, which perturbs both images and texts during AT. Following the theory of unimodal AT (Eq. 4), we formulate MAT objective as:
@@ 4.2.1 Multimodal inner maximization
Directly solving the inner-maximization in Eq. 6 is highly non-trivial due to the difficulty of updating both modalities simultaneously, and the increased computational cost.
To address these challenges, we emphasize that exact maximization is not always necessary; in fact, the basic PGD-AT defense method [madry2017towards] approximates maximization with first-order adversaries. Our extensive analysis led us to design MAT to be both effective and reasonably efficient, while also offering insights for future multimodal defenses.
Specifically, we mainly adopt two practical strategies: (1) Step-by-step perturbation: we generate adversarial examples $(I^{\prime},T^{\prime})$ by sequentially perturbing the text and image modalities, and (2) Loss simplification: replacing $\mathcal{L}_{\text{VL}}$ with effective yet efficient alternative.
Adversarial text generation. First, we generate adversarial texts using a representative text attack, BERT-attack [li2020bert]. BERT-attack identifies critical words contributing to the loss and replaces them with semantically similar, grammatically correct alternatives to maximize the loss. However, this discrete optimization process is computationally expensive, and directly maximizing $\mathcal{L}_{\text{VL}}$ over multiple image-text pairs in a batch by repeatedly replacing words would be prohibitively time-consuming. To address this, we exploit the shared image-text alignment objective in VL models (e.g., CLIP, ALBEF, BLIP), and instead of direct loss maximization, we maximize the divergence between individual image-text embeddings as:
$T^{\prime}=T+\operatorname*{arg\,max}_{\delta_{\text{T}}\in\Delta_{\text{T}}}-\frac{f_{\theta_{\text{I}}}(I)\cdot f_{\theta_{\text{T}}}(T+\delta_{T})}{\|f_{\theta_{\text{I}}}(I)\|\|f_{\theta_{\text{T}}}(T+\delta_{T})\|}.$ (7)
Adversarial image generation. Next, given adversarial texts $T^{\prime}$, we generate adversarial images $I^{\prime}$ using the widely adopted PGD attack [madry2017towards]. Unlike the discrete optimization in text attacks, image attacks involve continuous optimization, where each update step requires a single back-propagation to maximize a loss function. Therefore, directly maximizing the loss $\mathcal{L}_{\text{VL}}$ is practically feasible. For CLIP, instead of maximizing $\mathcal{L}_{\text{CLIP}}$, we adopt a simple approach by minimizing the cosine similarity between the image-text embeddings as:
$I^{\prime}=I+\operatorname*{arg\,max}_{\delta_{\text{I}}\in\Delta_{\text{I}}}-\frac{f_{\theta_{\text{I}}}(I+\delta_{\text{I}})\cdot f_{\theta_{\text{T}}}(T^{\prime})}{\|f_{\theta_{\text{I}}}(I+\delta_{\text{I}})\|\|f_{\theta_{\text{T}}}(T^{\prime})\|}.$ (8)